What is the radius of the nucleus of 28/14 Si ? What is its volume?



1
Expert's answer
2022-05-17T16:04:56-0400
"R={(1.3\:\rm fm)}\:A^{1/3}\\
=(1.3\:\rm fm)\:28^{1/3}=3.95\:\rm fm"

"V=\frac{4}{3}\pi R^3\\
=\frac{4}{3}*3.14*(\rm 3.95\: fm)^3=258\: fm^3"